OpenStudy (anonymous):

Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. (4 points and these are my equations y=(1.2)^x ( Neighborhood A) y=40+3x ( Neighborhood B) now when is it the same?

OpenStudy (astrophysics):

An easy way to do this would be to plug in the number of years in x, and see if they gave you the same number

OpenStudy (astrophysics):

Your first equation should be \[y=30(1.2)^x\]

OpenStudy (anonymous):

i tried doing that but it didnt work out as planned

OpenStudy (astrophysics):

Yes it maybe hard to do without a graph, so here is the graph, the point where it's the same is the intersection  :)

OpenStudy (anonymous):

how do u blow it up cause i cant see nothing

OpenStudy (astrophysics):

Hey, just go on this site and plug in the equation as I have on the left, I think if you see for yourself it will be much clearer https://www.desmos.com/calculator

OpenStudy (anonymous):

okay thx

OpenStudy (astrophysics):

Np

OpenStudy (nincompoop):

equate both of the functions and solve for zeroes

OpenStudy (nincompoop):

intersection

OpenStudy (astrophysics):

Naw, no need

OpenStudy (astrophysics):

Just use graph

OpenStudy (nincompoop):

is this calculus or just algebra?

OpenStudy (astrophysics):

algebra

OpenStudy (nincompoop):

sorry, my bad

OpenStudy (astrophysics):

It would be annoying to equate them as well haha

OpenStudy (nincompoop):

in case you need to test your algebra skills on logarithmic and and linear equations, this would be a good problem to solve for intersection (value of x and y)

OpenStudy (astrophysics):

Actually it's not too bad, would require you to know your exponent rules though

OpenStudy (astrophysics):

\[30\left( \frac{ 2^2 \times 3 }{ 2 \times 5 } \right)^2 = 40+x...\] etc

OpenStudy (anonymous):

this is not working out @astrophycsics

OpenStudy (anonymous):

thx to the person that gave me the metal

OpenStudy (astrophysics):

Np, don't let the negative value confuse you!